OpenStudy (anonymous):

A bowl of soup cools according to Newton's Law of Cooling. Its temperature (in *F) at the time of (t) is given bt T(t)=68+144e^-0.04t, where (t) is gven in minutes. Part I: What was the initial temperature (at the time of t=0) of the soup? Part II: What is the temperaure of the soup when t=15 minutes? Part III: To the nearest whole minute at what time (t) is the soups temerature 125 degrees Fahrenheit?

OpenStudy (zehanz):

So are given the formula that describes the temperature T as a function of the time t. All you need to do is substituting t=0 and t=15 for Parts I and II. For Part II, you have to solve T(t)=125 for t. Just try it!

OpenStudy (anonymous):

I did and I completely failed... Could you possibly give me a step by step instruction?

OpenStudy (zehanz):

Part I: set t=0: T(0)=68+144e^0. What is e^0?

OpenStudy (anonymous):

1?

OpenStudy (zehanz):

Yes, so T(0)=?

OpenStudy (anonymous):

212?

OpenStudy (zehanz):

True, again! See, it's not that hard! Now calculate T(15).

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Now i have another question how do i solve e^-0.04(15)?

OpenStudy (zehanz):

There is not much to solve here: -0.04*15=0.6, so all you need to do is grab a calculator and type e^0.6

OpenStudy (zehanz):

If you multiply that answer by 144 and then add 68, you've found T(15)!

OpenStudy (anonymous):

okay thanks:) i am going to have another question will you help me with it too?

OpenStudy (anonymous):

how do i solve part 3?

OpenStudy (zehanz):

Part 3: Solve T(t)=125 for t, so first write down the equation: \[68+144e^{-0.04t}=125\] Remember, you are looking for t, but it is still "up there" somewhere, so we must take a few steps to get there. First step?

OpenStudy (anonymous):

sutract 68?

OpenStudy (zehanz):

Go ahead! And then?

OpenStudy (anonymous):

divide by 144

OpenStudy (zehanz):

Do it! and then?

OpenStudy (anonymous):

Thats where im stuck

OpenStudy (zehanz):

OK, so you are here now:\[e^{-0.04t}=\frac{ 57 }{ 144 }\]:You can replace this immediately by:\[-0.04t=\ln \frac{ 57 }{ 144 }\]Why? Because of the definition of the logarithm, which says:\[\log_{a}b=c \Leftrightarrow a^c=b \]In this case, having e as base:\[lna=c \Leftrightarrow e^c=a\] So whenever you are looking for an exponent (which you are here), you need to come up with a logarithm. Remember: logarithms are exponents! E.G: solve \[10^t=25 \Leftrightarrow t=\log_{10}25 =\log 25\] So go on with \[-0.04t=\ln \frac{ 57 }{ 144 }\]now!

OpenStudy (anonymous):

0.396=e^-0.04?

OpenStudy (zehanz):

No, we already had -0.04t=ln(57/144)=-0.92676, so t=-0.92676/-0.04=23.2 min

OpenStudy (zehanz):

Just use your calculator fo find ln(57/144) and divide that answer by -0.04

OpenStudy (anonymous):

alright. i get it now

OpenStudy (anonymous):

\[1000=\frac{ 10000 }{ 5+1245e ^{-0.97t} }\] how do i solve this?

OpenStudy (anonymous):

hello?

OpenStudy (zehanz):

I'm back ;) Divide both sides by 1000 to make it more manageable...